OpenStudy (anonymous):

Find the derivative of f(x) = cos x, a = pi/2

OpenStudy (anonymous):

@zepdrix i ddi it but can u check my answer?

zepdrix (zepdrix):

sure :O what are you suppose to do? Just find the derivative then plug \(\large a\) in? Or does this have something to do with approximation?

OpenStudy (anonymous):

okay my teacher said there is an equation, so here is what i did f(x) = cos (x) f'(x) = -sin(x) x = a = pi/2 so f'(pi/2) = -sin (pi/2) = -1

zepdrix (zepdrix):

Your teacher said there is an equation? :) That's not much detail lolol Your derivative looks correct at least.

OpenStudy (anonymous):

yea she said try again there is an equation

OpenStudy (anonymous):

let me show you how i wrote it first okay?

OpenStudy (anonymous):

i mean the answer i wrote first

zepdrix (zepdrix):

k

OpenStudy (anonymous):

d/dx ( cos (x)) = -sin(x) -sin (pi/2) = -1, so the slope is -1, y = mx + b --> y = =1x + b cos (pi/2) = 0 so the coordinates pare are (pi/2, 0) so y = 0x + b, 0(pi/2) + b equals to 0, 0 = y = 0x + 0 = y = 0 for my final answer.

zepdrix (zepdrix):

Oh so like the previous problems, you're looking for an equation of the tangent line?

OpenStudy (anonymous):

yes sorry i made a mistake

OpenStudy (anonymous):

yes yes the equation of the tangent line

zepdrix (zepdrix):

\[\large y=-x+b\] From here you plugged in your point \(\large \left(\pi/2,\;0\right)\). \[\large 0=-\pi/2+b \qquad \qquad \rightarrow \qquad \qquad b=\pi/2\] I think you plugged your \(\large 0\) into the wrong spot.

OpenStudy (anonymous):

ohhh okay

OpenStudy (anonymous):

so instead of 0 = 0(pi/2) + b what would i write

zepdrix (zepdrix):

\[\huge (\color{royalblue}{\pi/2},\;\color{orangered}{0}) \qquad = \qquad (\color{royalblue}{x},\;\color{orangered}{y})\] Plug them into here, match up the colors.\[\huge \color{orangered}{y}=(-1)\color{royalblue}{x}+b\] I'm not sure why you're plugging a \(\large 0\) in for your -1....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so 0 = (-1)(pi/2) + b ?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thx

zepdrix (zepdrix):

np